# Patch-MLP-Based Predictive Control: Simulation of Upstream Pointing Stabilization for PHELIX Laser System

### v1

@@ Abstract
High-energy laser facilities such as PHELIX at GSI require excellent beam pointing stability for reproducability and relative independence for future experiments. Beam pointing stability has been traditionally achieved using simple proportional–integral–derivative (PID) control which removes the problem of slow drift, but is limited because of the time delay in knowing the diagnosis and the inertia in the mechanical system associated with mirrors. In this work, we introduce a predictive control strategy where the forecasting of beam pointing errors is performed by a patch-based multilayer perceptron (Patch-MLP) designed to capture local temporal patterns for more robust short-term jitter prediction. The subsequent conversion of these predicted errors into correction signals is handled by a PID controller. The neural network has been trained on diagnostic time-series data to predict beam pointing error. Using the feed-forward controller compensates for system delays. Simulations with a correction mirror placed upstream of the PHELIX pre-amplifier bridge confirm that the predictive control scheme reduces residual jitter compared to conventional PID control. Over a 10-hour dataset the controller maintained stable performance without drift, while standard pointing metrics showed consistent improvements of the order of $10\%\text{--}20\%$. The predictive controller operates without drift, and therefore may improve reproducibility and operational efficiency in high energy, low repetition rate laser experiment conditions.
@@ 2.1 Workflow of the simulation
Figure 2: Workflow of the Hybrid Feed-Forward MLP-Driven PID and Traditional PID Control. The inset shows the step-response of the steering mirror model used in our simulation.
As illustrated in Fig. 2, we present a hybrid feed-forward control scheme that combines an MLP time-series forecaster with a traditional PID controller to realize predictive beam pointing stabilization. The predicted positions are then converted into feed-forward commands for a tip–tilt mirror at the pre-Amplifier bridge.
As shown in Fig. 2, we use a second-order damped transfer function to represent the piezo-actuated tip-tilt mirror. The dynamics of the piezo-actuated tip-tilt mirror are approximated by a second-order mass–spring–damper system, as the relatively large mirror (e.g. 4 inch) mass dominates the response. This results in a single resonance mode that captures the essential dynamics, while higher-order modes of the piezo actuator are considered negligible within the operating frequency range. This lets us describe how its mechanical behavior affects the correction performance during simulation. The transfer function in Laplace spaces given by $G(s)=\omega_{n}^{2}/(s^{2}+2\zeta\omega_{n}s+\omega_{n}^{2}$). $\omega_{n}=2\pi f$ denotes the natural angular frequency of the mirror, where $f=200~\mathrm{Hz}$, and $\zeta=0.35$ is the damping ratio. The parameter values were chosen as a pragmatic approximation for a large 4 inch tip/tilt mirror, based on extrapolations from the PI S-340 datasheet towards heavier mirrors, sufficient for a demonstration model [17].
@@ 2.2 Patch MLP
Machine learning (ML) is a time-series forecasting method to predict the future behavior of a dynamical system. It has been applied in diverse fields such as finance, climate modeling, and industrial process control. Recently, it has been demonstrated that deep learning models can accurately predict the center position of a laser beam. These studies [16, 19] imply that the prediction time only needs to cover the model computation time and the optical system delay (dominated by the latency of the driver) [18]. Therefore, aiming for short prediction intervals, we employed a relatively lightweight architecture—an MLP—which provides sufficient predictive capability while maintaining low computational time.
@@
Figure 3: Patch-based MLP Architecture for Two-Axis Beam Pointing Prediction
@@
where $\theta$ denotes all trainable parameters. We use time-series data acquired by the PHELIX MAS [1] with a sampling interval of $T_{s}=0.244~\mathrm{ms}$. To account for both the per-step inference time of the model and the system latency, we set the prediction offset to $\Delta=5$ points, i.e., a horizon of $\Delta T_{s}\approx 1.22~\mathrm{ms}$, chosen according to $\Delta>\left\lceil(T_{\text{model}}+T_{\text{sys}})/T_{s}\right\rceil$. Predictions at offset $\Delta$ are then used for feed-forward compensation.
@@ 3 Simulation Results and Discussion
Before presenting the results, we briefly summarize the training procedure. The Patch-MLP forecaster, with about $2\times 10^{6}$ trainable parameters, was trained on one-minute time-series data ($\sim 2.4\times 10^{5}$ samples) for 50 epochs. Training on two NVIDIA V100 GPUs required about 3 minutes of wall-clock time.
@@ 3.1 Comparison of simulation results after compensation
Simulation results, presented in Fig. 4. The blue trace shows the uncontrolled signal; the orange trace is the output under a traditional PID controller with a $0.6\,\mathrm{ms}$ delay time; the green trace is the proposed MLP-driven PID that uses a prediction offset of $\Delta=5$ points ($\approx 1.22\,\mathrm{ms}$) to apply zero-delay feed-forward corrections. Both controllers markedly reduce the amplitude envelope compared with the real signal. And the proposed MLP-driven PID controller has better performance throughout the entire one-minute data.
Figure 4: One-minute beam-center trajectories (X and Y): real signal vs. traditional PID (0.6 ms delay) vs. MLP-driven PID.
Fig. 5(a) shows the improvement in laser pointing stability brought about by the new method. Under the control of the new method, the distribution of the laser beam center point is more concentrated in both the x-axis and the y-axis compared to the traditional PID method.
The MLP-driven PID (green) achieves a clear advantage below $60,\mathrm{Hz}$, where its power spectral density (PSD) remains consistently lower than that of the traditional PID (orange) on both axes (Fig. 5(b)). This indicates stronger suppression of slow drift and large-scale jitter components. The improvement arises from the prediction offset ($\Delta=5$ samples $\approx 1.22$,ms), which compensates for loop and mirror delays and reduces the phase lag that constrains a feedback-only PID in this frequency band. Consequently, while the conventional loop cannot fully attenuate the sub-$60$,Hz content, the predictive method achieves markedly lower residual amplitudes.
@@
Figure 5: Comparison of beam stabilization: (a) beam-center distribution and (b) power spectral density.
@@ 3.2 Model stability
For the long-term simulation, a single model was trained on one 1 min MAS segment and then kept fixed. The model was applied to all subsequent data collected over $\sim 10$ h. Acquisition was intermittent rather than strictly periodic: most intervals were 5 min, with occasional 10 min gaps. In total, 110 valid 1 min segments were obtained and used for evaluation. For each segment, predictions were generated and converted into simulated corrections following the procedure described above.
Figure 6: Average percentage reduction of residuals (MLP-driven PID vs. traditional PID) over the 10 h test. Positive values indicate lower residuals with the MLP-driven controller.
@@
Figure 7: Fourier comparison of failure segments and training data
@@
Overall, the predictive controller yields lower variance and a narrower distribution of the beam center on both axes, with no signs of drift across $10\text{\,}\mathrm{h}$.
@@ 4 Conclusion
Figure 8: Allan deviation analysis of the beam-pointing time series (1 min)

### v2

@@ Abstract
High-energy laser facilities such as PHELIX at GSI require excellent beam pointing stability for reproducability and relative independence for future experiments. Beam pointing stability has been traditionally achieved using simple proportional–integral–derivative (PID) control which removes the problem of slow drift, but is limited because of the time delay in knowing the diagnosis and the inertia in the mechanical system associated with mirrors. In this work, we introduce a predictive control strategy where the forecasting of beam pointing errors is performed by a patch-based multilayer perceptron (Patch-MLP) designed to capture local temporal patterns for more robust short-term jitter prediction. The subsequent conversion of these predicted errors into correction signals is handled by a PID controller. The neural network has been trained on diagnostic time-series data to predict beam pointing error. Using the hybrid predictive–feedback controller compensates for system delays. Simulations with a correction mirror placed upstream of the PHELIX pre-amplifier bridge confirm that the predictive control scheme reduces residual jitter compared to conventional PID control. In our simulations, over a 10-hour dataset the controller maintained stable performance without drift, while standard pointing metrics showed consistent improvements of the order of $10\%\text{--}20\%$. The predictive controller operates without drift, and therefore may improve reproducibility and operational efficiency in high energy, low repetition rate laser experiment conditions.
@@ 2.1 Workflow of the simulation
Figure 2: Workflow of the Hybrid predictive–feedback MLP-Driven PID and Traditional PID Control. The inset shows the step-response of the steering mirror model used in our simulation.
As illustrated in Fig. 2, we present a hybrid predictive–feedback control scheme that combines an MLP time-series forecaster with a traditional PID controller to realize predictive beam pointing stabilization. The predicted positions are then converted into commands for a tip–tilt mirror at the pre-Amplifier bridge.
As shown in Fig. 2, we use a second-order damped transfer function to represent the piezo-actuated tip-tilt mirror. The dynamics of the piezo-actuated tip-tilt mirror are approximated by a second-order mass–spring–damper system, as the relatively large mirror (e.g. 4 inch) mass dominates the response. This results in a single resonance mode that captures the essential dynamics, while higher-order modes of the piezo actuator are considered negligible within the operating frequency range. This lets us describe how its mechanical behavior affects the correction performance during simulation. The transfer function in Laplace space given by $G(s)=\omega_{n}^{2}/(s^{2}+2\zeta\omega_{n}s+\omega_{n}^{2}$). $\omega_{n}=2\pi f$ denotes the natural angular frequency of the mirror, where $f=200~\mathrm{Hz}$, and $\zeta=0.35$ is the damping ratio. The parameter values were chosen as a pragmatic approximation for a large 4 inch tip/tilt mirror, based on extrapolations from the PI S-340 datasheet towards heavier mirrors, sufficient for a demonstration model [18].
@@ 2.2 Patch MLP
Machine learning (ML) is a time-series forecasting method to predict the future behavior of a dynamical system. It has been applied in diverse fields such as finance, climate modeling, and industrial process control. Recently, it has been demonstrated that deep learning models can accurately predict the center position of a laser beam. These studies [2, 4] imply that the prediction time only needs to cover the model computation time and the optical system delay (dominated by the latency of the driver) [16]. Therefore, aiming for short prediction intervals, we employed a relatively lightweight architecture—an MLP—which provides sufficient predictive capability while maintaining low computational time. Transformer-based architectures are avoided because they are computationally more expensive, and recent studies show that their permutation-invariant self-attention fails to reliably capture temporal order in physical time-series, while simpler models often achieve equal or better short-horizon performance [21].
@@
Figure 3: Patch-based MLP Architecture for Two-Axis Beam Pointing Prediction.
@@
where $\theta$ denotes all trainable parameters. We use time-series data acquired by the PHELIX MAS [13] with a sampling interval of $T_{s}=0.244~\mathrm{ms}$ (sampling period). To account for both the per-step inference time of the model and the system latency, we set the prediction offset to $\Delta=5$ points, i.e., a horizon of $\Delta T_{s}\approx 1.22~\mathrm{ms}$, chosen according to $\Delta>\left\lceil(T_{\text{model}}+T_{\text{sys}})/T_{s}\right\rceil$. Predictions at offset $\Delta$ are then used for compensation.
@@ 3 Simulation Results and Discussion
Before presenting the results, we briefly summarize the training procedure. The Patch-MLP forecaster, with about $2\times 10^{6}$ trainable parameters, was trained on one minute of time-series data sampled at $T_{s}=0.244$ ms (yielding $\sim 2.4\times 10^{5}$ samples). The first $80\%$ of this sequence was used for training and the remaining $20\%$ for validation, ensuring a temporally consistent split without information leakage. The model was trained for 50 epochs, which took about three minutes of wall-clock time on two NVIDIA V100 GPUs.
@@ 3.1 Comparison of simulation results after compensation
Simulation results, presented in Fig. 4. The blue trace shows the uncontrolled signal; the orange trace is the output under a traditional PID controller with a $0.6\,\mathrm{ms}$ delay time; the green trace is the proposed MLP-driven PID that uses a prediction offset of $\Delta=5$ points ($\approx 1.22\,\mathrm{ms}$) to apply zero-delay corrections. Both controllers markedly reduce the amplitude envelope compared with the real signal. And the proposed MLP-driven PID controller has better performance throughout the entire one-minute data.
Figure 4: One-minute beam-center trajectories (X and Y): real signal vs. simulated traditional PID (0.6 ms delay) vs. simulated MLP-driven PID.
Fig. 5(a) shows the improvement in laser pointing stability brought about by the new method on the same one-minute dataset as used in Fig. 4, recorded two hours and 15 minutes after the training data. Under the control of the new method, the distribution of the laser beam center point is more concentrated in both the x-axis and the y-axis compared to the traditional PID method.
The MLP-driven PID (green) achieves a clear advantage below $60$ Hz, not through a uniform reduction over the entire band, but through a pronounced suppression of the dominant residual peaks that remain under conventional PID control on both axes (Fig. 5(b)). The improvement arises from the prediction offset ($\Delta=5$ samples $\approx 1.22$ ms), which compensates for loop and mirror delays and reduces the phase lag that constrains a feedback-only PID in this frequency band. Consequently, while the conventional loop cannot fully attenuate the sub-$60$ Hz content, the predictive method achieves markedly lower residual amplitudes. A comparison of cumulative RMS of feedback-only PID and MLP-driven PID can be found in the supplementary data.
@@
Figure 5: Comparison of simulated beam stabilization: (a) beam-center distribution and (b) power spectral density.
@@ 3.2 Model stability
For the long-term simulation, a single model was trained on one 1 min MAS segment and then kept fixed. The model was applied to all subsequent data collected over $\sim 10$ h. Acquisition was intermittent rather than strictly periodic: most intervals were 5 min, with occasional 10 min gaps. In total, 110 valid 1 min segments were obtained and used for evaluation. For each segment, predictions were generated and converted into simulated corrections following the procedure described above. The one-minute training segment and the 10 h test dataset are fully disjoint, with the latter taken after the training data on the same day. No test-set information was used for model training, and no hyperparameter tuning was performed on the test data. All hyperparameters were fixed before evaluation, and the trained model was applied to the 10 h dataset without further adjustment.
Figure 6: Average percentage reduction of residuals (MLP-driven PID vs. traditional PID) over the 10 h simulated test. Positive values indicate lower residuals with the MLP-driven controller.
@@
Figure 7: Fourier comparison of failure segments and training data.
@@
Overall, the predictive controller yields lower variance and a narrower distribution of the beam center on both axes, with no signs of drift across $10\text{\,}\mathrm{h}$. Because the model can be retrained within only a few minutes using approximately one minute of newly acquired data, it would also be straightforward to update the model should longer-term drift occur. In future work, the prediction accuracy could be further improved by distinguishing different classes of environmental vibration sources based on downstream experimental requirements, allowing the model to focus on the most relevant disturbance components.
@@ 4 Conclusion
The feasibility of deploying this method in a real-time system appears fundamentally sound, provided that several practical aspects are addressed. The required fast steering mirror is standard control hardware and can be procured off the shelf. The only missing component is an experimental characterization of a steering-mirror assembly of the necessary size. Our second-order actuator model, extrapolated from datasheet specifications (mirror mass and drive force), suggests comparable dynamic timescales once such a unit is characterized. The latency budget assumed in this study follows prior AO work at Apollon [16], where high-speed wavefront sensing and a real-time computer (RTC) operated at the relevant bandwidths; a similar platform already exists at PHELIX and can be extended in both software and hardware, with the sole structural addition being the inference latency, currently estimated at $\sim 0.3$ ms, which combined with the $\sim 0.6$ ms system delay remains well within the prediction capabilities of the model. A more fundamental integration challenge arises from the fact that the model assumes direct access to the raw perturbed tip/tilt signal, whereas in most laser systems the diagnostic sensor is located downstream of the steering mirror. This can be addressed either through an open-loop scheme—placing the sensor upstream of the actuator but downstream of the jitter source, supported by an accurate dynamical model of the mirror and optionally a secondary sensor for tuning—or through a pseudo-open-loop architecture that reconstructs the undisturbed input from corrected measurements and historical mirror commands using the actuator model, the latter offering greater generality at the cost of increased control-design complexity. Finally, an experimental evaluation path is available through the ALADIN project (funded by the German BMFRT), which is establishing the LAOTSE beam-control test bench at PHELIX, providing a suitable environment for realistic real-time performance assessment once the required hardware is in place.
@@
Figure 8: Allan Deviation Comparison: Real Data vs. Simulated PID and Simulated MLP-Driven PID (1 min).
Figure 9: Power spectral density of beam jitter (X/Y, 0–300 Hz).
Figure 10: Open-Loop Bode Plot: Mirror and PID Controller.
Figure 11: Error in predicted value calculation vs. Error caused by a 0.6ms delay.
Figure 12: Power spectral density of beam jitter (without simulation).
Figure 13: Cumulative RMS of Residual Jitter: Simulated Traditional PID and Simulated MLP-Driven PID.